For a recent 2-year period, the balance sheet of Wildhorse Company showed the following stockholders’ equity data at December 31 (in millions). 2020 2019 Additional paid-in capital $ 900 $ 831 Common stock 651 642 Retained earnings 7,210 5,240 Treasury stock 1,755 1,008 Total stockholders’ equity $7,006 $5,705 Common stock shares issued 217 214 Common stock shares authorized 500 500 Treasury stock shares 39 28 (a) Answer the following questions. (1) What is the par value of the common stock? (Round par value to 2 decimal places, e.g. $3.15.) Par value of common stock $enter the par value of the common stock rounded to 2 decimal places (2) What is the cost per share of treasury stock at December 31, 2020, and at December 31, 2019? December 31, 2020 December 31, 2019 Cost per share of Treasury stock $enter a dollar amount $enter a dollar amount (b) Prepare the stockholders’ equity section at December 31, 2020. (Enter account name only and do not provide descriptive information.)
Answer #1

Answer:

1) Par value per common share = 651/217 = $3 per share
b) Average cost per share
December 31,2020 December 31,2019
Cost per share 1755/39 = 45 1008/28 = 36
c) Stockholder's equity
Stockholder's equity Amount($)
Paid in capital
Common Stock                                     651
Additional paid in capital                                     900
Total paid in capital                                  1,551
Retained earnings                                  7,210
Total                                  8,761
Less:
Treasury stock                                -1,755
Total Stockholder's equity                                  7,006